change each percent to a fraction. simplify when necessary. 2/3%
Answer by stanbon(75887)

change each percent to a fraction. simplify when necessary. 2/3%
---------------
(2/3)% = (2/3)/100 = 0.66666666.../100 = 0.0066666666...
= 66.6666666../10,000
Cheers,
Stan H.